OpenStudy (jrmeyer):

If the system y=x^2 + 2x + 3, y = kx-1, has exactly one solution, then what does k equal?

OpenStudy (cwrw238):

for one solution b^2 - 4ac = 0 ( for ax^2 + bx + c = 0) so here kx - 1 = x^2 + 2x + 3 x^2 + 2x - kx + 4 = 0 x^2 + (2 - k)x + 4 = 0 b^2 - 4ac = (2-k)^2 - 4*4 = 0 solve this for k

OpenStudy (cwrw238):

can u solve this last equation?

OpenStudy (jrmeyer):

no - not following

OpenStudy (cwrw238):

(2-k)^2 = 16 take square root of both sides 2-k = 4 or -4 so k = 6 or -2

OpenStudy (cwrw238):

what didnt you follow the b^2 - 4ac bit??

OpenStudy (jrmeyer):

I thought in a system question, it was one or the other. I think the part where you had them equal each other was throwing me off.

OpenStudy (cwrw238):

thats called the discriminant and its value indicates how many real roots there are if its = 0 there is one real root,if postive there are 2 real roots and if its negative there are no real roots

OpenStudy (cwrw238):

the two expression k - 1 and x^2 =2x + 3 are BOTHe qual to y therefore they must be equal to each other.

OpenStudy (jrmeyer):

got it - thanks!

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

@cwrw238 but there can only be one answer?